OpenStudy (anonymous):

Find w(-3) using synthetic substitution: w(x)=11x^3-6x^2+2 I got -241, can someone confirm if this is correct?

OpenStudy (tkhunny):

Don't think so. Did you use the weird, missing 2rd term, 0x?

OpenStudy (anonymous):

Oh, i dont think so! I did: w(-3)=11(-3)^3-6(-3)^2+2 11(-27)-6(-9)+2 -297+54+2 -297+56=-241 @tkhunny

OpenStudy (anonymous):

so how would I do it with what you said?

OpenStudy (anonymous):

11(-3)^3-6(-3)^2+2 11(-27)-6(-9)+2 can you see how the -9 is close but a bit off? (-3)^2 = (-3)*(-3)= ?

OpenStudy (anonymous):

OH it should have been positive 9!

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

so then really it should be 11(-27)-6(9)+2? @DemolisionWolf

OpenStudy (anonymous):

yep! @ali1029

OpenStudy (anonymous):

so it would be -349?

OpenStudy (anonymous):

you got it ^_^

OpenStudy (anonymous):

thank you so much :) where would I be without you?!

OpenStudy (anonymous):

haha, some one would show up and just tell you the Answer, i'm sure of it

OpenStudy (anonymous):

haha, well it is definitely better being told how to solve the problem and getting my answer confirmed :p

OpenStudy (tkhunny):

That's just substitution, not "synthetic substitution".